# Adversarial Robustness for Deep Learning-based Wildfire Prediction Models

### v2

@@ 3 Wildfire Adversarial Robustness Procedure
In this section, we introduce the proposed framework, WARP (Wildfire Adversarial Robustness Procedure), for comprehensively assessing the robustness of smoke detection models. WARP generates two types of adversarial examples through data augmentation: global noise (i.e., Gaussian noise) and local noise (i.e., cloud PNG patches). We then compare the original detection (red bounding box) with the perturbed detection (blue bounding box): we conduct a sanity check for images perturbed global noise, which analyzes the effect of global noise on model precision (see Section 3.1). Additionally, we conduct the deception test for images perturbed local noise, which analyzes the robustness of the model’s localization and classification abilities on specific objects (see Section 3.2). We later demonstrate in Section 5 that the tests outlined above can identify vulnerabilities in robustness for wildfire detection models, which can produce data augmentation strategies tailored specifically for improving robustness for DNN models.
Figure 2: WARP workflow.
@@ 3.1 Global Sanity Check: Noise Overlay
For a given noise level, the mAP (mAP50-95) Percentage Loss $\mathcal{L}$ is calculated to quantify the precision loss from the addition of random noise. We choose mAP (see A.1 for detailed definitions) as the primary challenge metric since fixed threshold-precision metrics (i.e., mAP33, mAP50, mAP75, etc.) may only offer a limited evaluation of precision. mAP measures the mean average precision across Intersection over Union (IoU) thresholds ranging from 0.50 to 0.95, making it a robust metric [2].
@@ 3.2 Local Deception Test: Noise Patch
The previous test considers only the context of noise level variability via $\sigma$. We also proposes the local deception test, which introduces spatial context specific to the smoke detection task. Specifically, local noise patches (small images in the PNG (portable network graphics) format) are used to check for robustness against localized perturbations specific to wildfires. For each image $i$, a PNG patch of constant size is injected at a specific spot in the image (see Figure 3 (a) for instance). For computational efficiency, we divide each image into 25 by 25 grids, and the noise is injected in the center of each grid slot. “Noise” can be any wildfire-related object, including smoke-like objects such as clouds, or other objects in context, such as trees, buildings, glare, etc. We specifically used clouds as they are the most common subject of false positives in previous works [7, 10, 2]. Figure 3 (b) shows the cloud PNG used in the local deception test. Despite being horizontally wider than wildfire smoke, they represent the everyday cumulus-type clouds, which is why we selected it. After observing existing wildfire smoke data, and testing different configurations, we configured the patch to be $25\times 25$ pixels at 100% brightness.
To quantify the robustness against patch noise, three metrics are proposed, $\alpha_{i},\beta_{i}$, and $\gamma_{i}$. They are defined in Eqs. (4), (5), and (6) in the following sections.
@@ 3.2.1 Classification Flip Probabilities
Wildfire smoke detection models’ performance can be divided into classification and localization abilities. The classification ability refers to how precisely the model identifies an image as smoke-positive or smoke-negative. The localization ability refers to how precisely the model locates the smoke with bounding boxes given a smoke-positive image.
The Classification Flip Probabilities focuses on the vulnerability of the model’s classification abilities. When smoke-positive input images are given, a model may fail to generate bounding boxes for some images. These misclassified images can be called false negatives (FN). On the other hand, if the model successfully generates at least one bounding box in any location of an image, the image can be called a true positive (TP).
The Classification Flip Probabilities quantify how many TPs or FNs are “flipped” upon injecting local noise. Unlike the global sanity check test, noise injection is done grid-wise. Image $i$ is divided into $A_{i}=25\times 25=625$ equal grids, and local noise is injected into each grid slot. The classification outcome is then observed $A_{i}$ times in total for the image. For each image $i$, depending on whether it is a TP or FN, the classification vulnerability of a model can be quantified by calculating two conditional probabilities:
@@
where $l_{\text{null}}$ and $l_{\text{smoke}}$ denote the null and smoke classes. $\mathbb{I}_{i,j}(\cdot)$ is the indicator function, which equals 1 if the specified flip occurs for image $i$ when the noise is injected into the $j$-th grid slot. It equals 0 otherwise. Again, $A_{i}$ is fixed at $25\times 25=625$ for every image.
These image-wise probabilities are averaged over the entire test set data to evaluate the model’s classification vulnerabilities. In summary, $\alpha_{i}$ and $\beta_{i}$ represents the prediction robustness that image $i$ is smoke-positive and smoke-negative, respectively.
@@ 3.2.2 Localization Deception Rate
where $D_{i}$ is the number of detections and its bounding box has $IoU\geq 0.50$ (see Equation 13 for further details) with the injected noise’s bounding box. $A_{i}$ is the number of attempts, or the number of possible positions for the noise to be injected in a $25\times 25$ grid ($25^{2}=625)$. Since $D_{i}$ is discrete, $\gamma_{i}$ is also discrete. That is, $\{1,2,3,\dots,625\}$ is the set of all possible values of $D_{i}$, and the set of all possible deception rates $\gamma_{i}$ is $\{1/625,2/625,3/625,\dots,625/625\}=\{0.0016,0.0032,0.0048,\dots,1.0000\}$.
@@ 3.3 Data Collection and Preparation
The NEMO dataset was originally created to fine-tune DETR. It contains 2,500+ images, with roughly 90% being smoke-positive. The smoke-positive regions in the original video frames were cropped and zoomed for efficient training. Following [2], we used this dataset for model training. In contrast, the dataset by [4] contains more realistic 1,661 unedited smoke-positive images. We used this dataset for model testing and robustness evaluation. Table 1 shows the data split.
@@ 3.4 Object Detection Framework
  • For a CNN-based real-time object detection model, we choose YOLOv8, the 8th generation model of the YOLO (You Only Look Once) framework. It is a popular real-time object detection frameworks and is publically available on the Ultralytics API [24].
@@
Pre-trained weights were transferred from COCO-dataset-trained versions of YOLOv8 and RT-DETR. Their lightweight versions, namely YOLOv8-nano (YOLOv8n) and RT-DETR-large (RT-DETR-l), were chosen to reduce computation time during robustness evaluation. Default training and inference hyperparameters were used from [24] except for those shown in Table 2. Extensive hyperparameter tuning is encouraged in the future.
We use metrics mAP and mAP50 to measure model precision (See A.1 for detailed definitions), which are common challenge metrics in object detection. All training, validation, and testing were run on a Python 3.8.12 virtual environment with CUDA version 12.2, equipped with quadruple NVIDIA GeForce GTX 1080 GPUs.
@@ 4.2 Results of Global Sanity Check
Figure 5 shows the result of the global sanity check as the noise level $a$ is increased in increments of 0.001 from 0.0 to 0.4 over 400 iterations. Interestingly, we observed a substantial difference between the two architectures. Specifically, the mAP Percentage Loss $\mathcal{L}$ for the transformer-based RT-DETR-l generally degraded at a higher rate than that of the CNN-based YOLOv8n. $\mathcal{L}$ for RT-DETR-l converged to $\approx$100 at noise level $a\approx 0.300$, whereas YOLOv8n did not show signs of convergence even after the terminal noise level 0.400. $\mathcal{L}_{\text{RT-DETR-l}}(0.400)=99.23$ and $\mathcal{L}_{\text{YOLOv8n}}(0.400)=89.53$.
@@
To the naked eye, the noise is barely visible, only appearing as a dimming in image luminosity. There was minimal difference between the performance of the two models without noise (compare Figure 7 (b) with Figure 7 (a)). However, it is important to note that both models suffered in precision. YOLOv8n tended to be more conservative with its detections, only detecting smoke with a definitive white color, or in other words, at a later period of the incipient stage (see Figure 7 (b)). On the other hand, RT-DETR-l did not have this bias but instead began to confuse the sky as smoke with relatively high confidence (see Figure 7 (b)).
Further observation shows that both models may confuse clouds with smoke under noise stress (compare Figure 9 (a) with Figure 9 (b)), but RT-DETR-l makes false positives even when no clouds exist (see Figure 9 (b)). Again, YOLOv8n became conservative with its detections, losing detection confidence and sometimes not detecting at all (see Figure 9 (a)). Figures 9 and 9 are two representative examples of false detections made by both models for identical images, again at noise level $a=0.210$.
@@ 4.3 Results of Local Deception Test
In this section, we summarize the results of the local deception test, namely the classification flip probabilities and the localization deception rate. They offer a nuanced review of model robustness and how it translates to detection results. In addition, we conduct an auxiliary test was conducted for further insight.
@@ 4.3.1 Result of Classification Flip Probabilities
The results are summarized in Table 4. As discussed, $\alpha_{i}$ shows the robustness of the prediction that image $i$ is smoke-positive, and $\beta_{i}$ shows the robustness of the prediction that image $i$ is smoke-negative. The results indicate a trend of wildfire smoke detection. Since smoke’s features are subtle, the smoke-positive prediction is easily flipped by injecting a cloud-like noise into even one of the 625 grids. This strongly suggests that there is substantial room for improvement in the model, possibly through data augmentation with additional images containing smoke-like objects, as discussed in Section 5. On the other hand, both models show strong robustness for smoke-free predictions. In particular, no smoke-to-null flips were observed for RT-DETR-l, suggesting a greater potential of the transformer-based model for general object detection tasks.
@@
Here, RT-DETR-l’s self-attention mechanism may work in its favor for local perturbation classification. The architecture may help defend against classification changes under local perturbations, as demonstrated in Table 4. This can be because of the self-attention mechanism in transformers, which captures and leverages global relationships across the image. Unlike Section 4.2, where the entire image was perturbed, local perturbations may not substantially affect the global relationships within the original image.
@@ 4.3.2 Results of Localized Deception Rate
Table 5 shows the count of each possible $\gamma_{i}$ value for YOLOv8n and RT-DETR-l. The occurrence of local deception is relatively rare in both models. We also calculated the expected deception rate:
@@
Contrary to Section 4.3.1, RT-DETR-l’s self-attention may work to its deficit in local perturbation localization. Section 3.4 demonstrated that RT-DETR’s transformer architecture may cause challenges in small-object detection. In particular, object queries representing smoke’s subtle features may be de-prioritized, leading to worse precision for small objects. This causes bounding box confusion with similar small smoke-like objects like the cloud PNG patch. This is consistent with Table 5, as RT-DETR-l frequently recorded higher $\gamma_{i}$ values, and its localization deception rate was more than three times that of YOLOv8n. However, more testing is required to determine the significance of these results in real conditions.
@@
For further analysis, we extracted RT-DETR-l detections with a $\gamma_{i}>0.0000$, and compared to their YOLOv8n counterpart. We highlight the region where the noise patch deceived the model. There were 5 images where both models detected with a $\gamma_{i}>0.0000$. Interestingly, 4/5 of these detections were taken from the same surveillance camera, namely from the SMER TCS9 Site. Figures 11 (a) and 11 (b) show two such cases. There were several cases where RT-DETR-l recorded multiple deceptions whereas YOLOv8n remained completely unaffected. In particular, there were cases in which RT-DETR-l detected with a $\gamma_{i}$ value of 0.0032 and 0.0048 (see Figures 11 (b) and 12 (b), respectively), but YOLOv8n remained unchanged and detected with $\gamma_{i}$ value of 0.0000 (see Figures 11 (a) and 12 (a)). Below are identical images detected by the two models, where each grid represents the position at which the cloud-noise was injected. Red areas indicate noise-affected regions, whereas blue regions indicate otherwise.
@@
Factoring out rare cases where the cloud noise completely overlapped the smoke (this could only happen in one or two grids, putting the probability at around 0.16-0.32%), both models were deceived when the cloud PNG patch was near the smoke. This is alarming, since the cameras are positioned at vantage points, and clouds are more clearly visible at higher altitudes. Thus, the occlusion of smoke by clouds would be a relatively common occurrence. Furthermore, it can be explained that the models confuse smoke with cloud since both are similar in features, and the parts of the cloud PNG patch merge with the smoke at times. As seen with YOLOv8n, the models tend to reliably detect when the smoke has developed a clear white coloration. Thus, the models may have interpreted the cloud as the origin point of the smoke, even when the direction of the smoke plumes suggested otherwise.
To test if there was a spatial relationship between the deceptions, we enumerated the cumulative number of noise-affected detections for each slot that the patch noise was injected for both models. Both models exhibited a slight bias to be deceived at the center regions (see Figures 13 (a) and 13 (b)). A potential explanation is that smoke is most frequently depicted in the middle horizontal of the image. It becomes easier for human observers to spot and annotate as rises above the horizon, which is frequently in this region. This creates an annotation bias in the test set (see Figure 14). Thus, cloud PNG patches may appear as part of the smoke in this zone.
@@ 5 Discussion
Based on our analysis, we propose the following data augmentation strategies. These solutions should provide future training data for first-generation YOLOv8 and RT-DETR-based wildfire detection models. While these preliminary solutions appear to be promising, a more specific evaluation of adversarial robustness for proceeding generation models should be a topic of future research. Because this framework is model-agnostic, even if new DNN architectures are incorporated into wildfire smoke detection, these solutions may still be applied to diversify the limited quantity of wildfire smoke data. We propose the following data augmentation techniques for robustification and their reason for implementation.
@@
    • Speed-optimized transformers have lower resilience to Gaussian noise compared to CNNs. Gaussian noise with $0.1\lesssim a\lesssim 0.4$ should be introduced into speed-optimized transformers’ training data to bring its basic sanity level on par with CNNs.
    • While CNNs shows more resilience to Gaussian noise than transformers, it is still not immune. Gaussian noise at different levels should be introduced into both architectures’ training data to increase detection confidence under noise stress.
@@
    • Speed-optimized transformers exhibit dangerously low robustness for localization when cloud PNG patches are injected. Given that most deceptions occur in the middle-horizontal areas of the image, patches should be put in this region to improve confidence when clouds and smoke overlap. The same precautions should be taken for CNNs.
    • Furthermore, transformers and CNNs alike mistake clouds as smoke under whole-image perturbation. Cloud PNG patches should also be inserted into the upper areas (areas depicting the sky) for images already injected with Gaussian noise.
@@
    • An effective solution to combat false positives extensively used by [2] is to use collages. This is because collages help to better differentiate between smoke-positive and smoke-negative images. Since CNNs suffered from local perturbation classification, it collages of images between classes should be implemented into training data.
    • Furthermore, YOLOv4-style mosaic [26] has the added benefit of introducing object size variety in the data. This is particularly useful for small object detection. Collages inevitably make the already-small smoke object even smaller. Especially for speed-optimized transformers, crucial smoke features must be extracted first. The below augmentation strategy seeks to offer a potential solution.
@@
    • Data collected by [4] from [19] exclusively depicts smoke that appears at or near the horizon (i.e., the middle-horizontal of the image). Cropping a subset of images into equal quadrants will add spatial diversity of the position of the smoke.
    • Furthermore, since the crops will result in a smaller-sized image, when resized to the target image size ($640\times 640$ pixels), smoke will appear larger than usual. This may help models better extract the subtle features of smoke, offering a solution to the problem discussed in the previous augmentation strategy.

### v3

@@ 3 Wildfire Adversarial Robustness Procedure
In this section, we introduce the proposed framework, WARP (Wildfire Adversarial Robustness Procedure), for comprehensively assessing the robustness of smoke detection models. WARP generates two types of adversarial examples through data augmentation: global noise (i.e., Gaussian noise) and local noise (i.e., cloud PNG patches). We then compare the original detection (red bounding box) with the perturbed detection (blue bounding box): we conduct a sanity check for the images’ perturbed with global noise, which analyzes the effect of global noise on model precision (see Section 3.1). Additionally, we conduct the deception test for images perturbed with local noise, which analyzes the robustness of the model’s localization and classification abilities on specific objects (see Section 3.2). We later demonstrate in Section 5 that the tests outlined above can identify vulnerabilities in robustness for wildfire detection models, which can produce data augmentation strategies tailored specifically for improving the robustness of DNN models.
@@ 3.1 Global Sanity Check: Noise Overlay
For a given noise level, the mAP (mAP50-95) percentage loss $\mathcal{L}$ is calculated to quantify the precision loss from the addition of random noise. We choose mAP (see Appendix A.1 for detailed definitions) as the primary challenge metric since fixed threshold-precision metrics (i.e., mAP33, mAP50, mAP75, etc.) may only offer a limited evaluation of precision. mAP measures the mean average precision across Intersection over Union (IoU) thresholds ranging from 0.50 to 0.95, making it a robust metric [39].
@@ 3.2 Local Deception Test: Noise Patch
The previous test considers only the context of noise level variability via $\sigma$. We also propose the local deception test, which introduces spatial context specific to the smoke detection task. Specifically, local noise patches (small images in the PNG (portable network graphics) format) are used to check for robustness against localized perturbations specific to wildfires. For each image $i$, a PNG patch of constant size is injected at a specific spot in the image (see Figure 3a for instance). For computational efficiency, we divide each image into 25 by 25 grids, and the noise is injected in the center of each grid slot. “Noise” can be any wildfire-related object, including smoke-like objects such as clouds, or other objects in context, such as trees, buildings, glare, etc. We specifically used clouds as they are the most common subject of false positives in previous works [22, 1, 39]. Figure 3b shows the cloud PNG used in the local deception test. Despite being horizontally wider than wildfire smoke, they represent the everyday cumulus-type clouds, which is why we selected them. After observing existing wildfire smoke data, and testing different configurations, we configured the patch to be $25\times 25$ pixels at 100% brightness.
To quantify the robustness against patch noise, three metrics are proposed, $\alpha_{i},\beta_{i}$, and $\gamma_{i}$. They are defined in Equations (4)–(6) in the following sections.
@@ 3.2.1 Classification Flip Probabilities
A wildfire smoke detection model’s ability can be subdivided into its classification and localization abilities. The classification ability refers to how precisely the model can identify an image as smoke-positive or smoke-negative. The localization ability refers to how precisely the model can locate the smoke object using bounding boxes given a smoke-positive image.
The Classification Flip Probabilities focuses on the vulnerability of the model’s classification abilities. When smoke-positive input images are given, a model may fail to generate bounding boxes for some images. These misclassified images can be called false negatives (FNs). On the other hand, if the model successfully generates at least one bounding box in any location of an image, the image can be called a true positive (TPs).
The Classification Flip Probabilities quantify how many TPs or FNs are “flipped” upon injecting local noise. Unlike the global sanity check test, noise injection is performed grid-wise. Image $i$ is divided into $A_{i}=25\times 25=625$ equal grids, and local noise is injected into each grid slot. The classification outcome is then observed $A_{i}$ times in total for the image. For each image $i$, depending on whether it is a TP or FN, the classification vulnerability of a model can be quantified by calculating two conditional probabilities:
@@
where $l_{\text{null}}$ and $l_{\text{smoke}}$ denote the null and smoke classes. $\mathbb{I}_{i,j}(\cdot)$ is the indicator function, which equals 1 if the specified flip occurs for image $i$ when the noise is injected into the $j$-th grid slot. It equals 0 otherwise. Again, $A_{i}$ is fixed at $25\times 25=625$ possible slots for every image.
These image-wise probabilities are averaged over the entire test set data to evaluate the model’s classification vulnerabilities. To summarize what these metrics mean, $\alpha_{i}$ and $\beta_{i}$ represent the classification robustness that image $i$ is smoke-positive and smoke-negative, respectively.
@@ 3.2.2 Localization Deception Rate
where $D_{i}$ is the number of detections and its bounding box has $IoU\geq 0.50$ (see Equation (13) for further details) with the injected noise’s bounding box. $A_{i}$ is the number of attempts, or the number of possible positions for the noise to be injected in a $25\times 25$ grid ($25^{2}=625)$. Since $D_{i}$ is discrete, $\gamma_{i}$ is also discrete. That is, $\{1,2,3,\dots,625\}$ is the set of all possible values of $D_{i}$, and the set of all possible deception rates $\gamma_{i}$ is $\{1/625,2/625,3/625,\dots,625/625\}=\{0.0016,0.0032,0.0048,\dots,1.0000\}$.
@@ 3.3 Data Collection and Preparation
The NEMO dataset was originally created to fine-tune DETR. It contains 2500+ images, with roughly 90% being smoke-positive. The smoke-positive regions in the original video frames were cropped and zoomed for efficient training. Following [39], we used this dataset for model training. In contrast, the dataset by [17] contains more realistic 1661 unedited smoke-positive images. We used this dataset for model testing and robustness evaluation. Table 1 shows the data split.
@@ 3.4 Object Detection Framework
  • For a CNN-based real-time object detection model, we choose YOLOv8, the 8th generation model of the YOLO (You Only Look Once) framework. It is a popular real-time object detection framework and is publically available on the Ultralytics API [35].
@@
Pre-trained weights were transferred from COCO-dataset-trained versions of YOLOv8 and RT-DETR (both v8.3.67). Their lightweight versions, YOLOv8-nano (YOLOv8n) and RT-DETR-large (RT-DETR-l), were chosen to reduce computation time during robustness evaluation. Default training and inference hyperparameters were used from [35]. Training YOLOv8n for 250 epochs and RT-DETR-l for 285 produced our best results. However, extensive hyperparameter tuning is encouraged in the future.
We use metrics mAP and mAP50 to measure model precision (See Appendix A.1 for detailed definitions), which are common challenge metrics in object detection. All training, validation, and testing were run on a Python 3.8.12 virtual environment with CUDA version 12.2, equipped with quadruple NVIDIA GeForce GTX 1080 GPUs from the University of Nevada, Reno.
@@ 4.2 Results of Global Sanity Check
Figure 5 shows the result of the global sanity check as the noise level $a$ is increased in increments of 0.001 from 0.0 to 0.4 over 400 iterations. Interestingly, we observed a substantial difference between the two architectures. Specifically, the mAP percentage loss $\mathcal{L}$ for the transformer-based RT-DETR-l generally degraded at a higher rate than that of the CNN-based YOLOv8n. $\mathcal{L}$ for RT-DETR-l converged to $\approx$100 at noise level $a\approx 0.300$, whereas YOLOv8n did not show signs of convergence even after the terminal noise level 0.400 ($\mathcal{L}_{\text{RT-DETR-l}}(0.400)=99.23$ and $\mathcal{L}_{\text{YOLOv8n}}(0.400)=89.53$).
@@
To the naked eye, the noise is barely visible, only appearing as a dimming in image luminosity. There was minimal difference between the performance of the two models without noise (compare Figure 7b with Figure 7a). However, it is important to note that both models suffered in precision. YOLOv8n tended to be more conservative with its detections, only detecting smoke with a definitive white color, or in other words, at a later period of the incipient stage (see Figure 7b). On the other hand, RT-DETR-l did not have this bias but instead began to confuse the sky for smoke with relatively high confidence (see Figure 7b).
Further observation shows that both models may confuse clouds with smoke under noise stress (compare Figure 9a with Figure 9b), but RT-DETR-l makes false positives even when no clouds exist (see Figure 9b). Again, YOLOv8n became conservative with its detections, losing detection confidence and sometimes not detecting at all (see Figure 9a). Figures 9 and 9 are two representative examples of false detections made by both models for identical images, again at noise level $a=0.210$.
@@ 4.3 Results of Local Deception Test
In this section, we summarize the results of the local deception test, which included an analysis of the classification flip probabilities and the localization deception rate. In addition, we conduct an auxiliary test that contextualizes the role of image annotations in our findings.
@@ 4.3.1 Result of Classification Flip Probabilities
The results are summarized in Table 4. As discussed, $\alpha_{i}$ shows the robustness of the prediction that image $i$ is smoke-positive, and $\beta_{i}$ shows the robustness of the prediction that image $i$ is smoke-negative. The results indicate a trend in wildfire smoke detection. Since smoke’s features are subtle, the smoke-positive prediction is easily flipped by injecting a cloud-like noise into even one of the 625 grids. This strongly suggests that there is substantial room for improvement in the model, possibly through data augmentation with additional images containing smoke-like objects, as discussed in Section 5. On the other hand, both models show strong robustness for smoke-free predictions. In particular, no smoke-to-null flips were observed for RT-DETR-l, suggesting a greater potential of the transformer-based model for general object detection tasks.
@@
Here, RT-DETR-l’s self-attention mechanism may work in its favor for local perturbation classification. The architecture may help defend against classification changes under local perturbations, as demonstrated in Table 4. This may be attributable to the self-attention mechanism in transformers, which leverages global relationships across the image. Unlike Section 4.2, where the entire image was perturbed, local perturbations may not substantially affect the global relationships.
@@ 4.3.2 Results of Localized Deception Rate
Table 5 shows the frequency of observed $\gamma_{i}$ values. We also calculated the expected deception rate.
@@
Contrary to Section 4.3.1, RT-DETR-l’s self-attention may work to its deficit in local perturbation localization. Section 3.4 demonstrated that RT-DETR’s transformer architecture may cause challenges in small-object detection. In particular, object queries representing smoke’s subtle features may be de-prioritized, leading to worse precision for small objects. This causes bounding box confusion with similar small smoke-like objects like the cloud PNG patch. This is consistent with Table 5, as RT-DETR-l frequently recorded higher $\gamma_{i}$ values, and its localization deception rate was more than 3 times that of YOLOv8n. However, more testing is required to determine the practical significance.
@@
For further analysis, we extracted RT-DETR-l detections where $\gamma_{i}>0.0000$, and compared them to their YOLOv8n counterpart. We highlighted where the noise patch deceived the model. There were five images where both models detected with a $\gamma_{i}>0.0000$. Interestingly, four out of five of these detections were taken from the SMER TCS9 surveillance site. Figure 11a,b show two such cases. There were several cases where RT-DETR-l recorded multiple deceptions whereas YOLOv8n remained completely unaffected. In particular, there were cases in which RT-DETR-l detected with a $\gamma_{i}$ value of 0.0032 and 0.0048 (see Figures 11b and 12b, respectively), but YOLOv8n remained unchanged and detected with $\gamma_{i}$ value of 0.0000 (see Figures 11a and 12a). Below are identical images detected by the two models, where each grid shows where the cloud-noise was injected. Red areas indicate noise-affected regions.
@@
Factoring out rare cases where the cloud noise completely overlapped the smoke (this could only happen in one or two grids, putting the probability at around 0.16–0.32%), both models were deceived when the cloud PNG patch was near the smoke. This is alarming, since the cameras are positioned at vantage points, and clouds are more clearly visible at higher altitudes. Thus, the occlusion of smoke by clouds would be a relatively common occurrence. Furthermore, it can be explained that the models confuse smoke with cloud since both are similar in features, and the parts of the cloud PNG patch merge with the smoke at times. As seen with YOLOv8n, the models tend to reliably detect when the smoke has developed a clear white coloration. Thus, the models may have interpreted the cloud as the origin point of the smoke, even when the direction of the smoke plumes suggested otherwise.
To test if there was a spatial relationship between the deceptions, we enumerated the cumulative number of noise-affected detections for each slot that the patch noise was injected for both models. Both models exhibited a slight bias to be deceived at the center regions (see Figure 13a,b). A potential explanation is that smoke is most frequently depicted in the middle horizontal of the image. It becomes easier for human observers to spot and annotate as smoke rises above the horizon, which is frequently in this region. This creates an annotation bias in the test set (see Figure 14), which may have caused the models to over-focus on this region. Thus, cloud PNG patches may appear as part of the smoke most in this zone.
@@ 5 Discussion
In a comprehensive comparative study, we analyze the adversarial robustness of CNN- and transformer-based real-time smoke detection models in context to other studies. The main findings are summarized below:
  • The global sanity check revealed that the transformer-based RT-DETR-l model was substantially more vulnerable to global noise injection compared to the CNN-based YOLOv8, even at noise levels barely visible to the human eye. This is a notable contrast from other studies [25, 31] which suggest that transformers trained on Big Data (i.e., CIFAR-10, MNIST, etc.) are more robust than CNNs against global noise. The current results may be an artifact of the severe data shortage in wildfire detection. Thus, as more data augmentation techniques (such as those from WARP) create more diverse adversarial examples, transformers will most likely overtake CNNs in terms of global noise robustness.
  • An analysis of the classification flip probabilities revealed that both CNN- and transformer-based models are sensitive to local noise injection. A single noise–injected grid (out of 625 total) resulted in flip probabilities of $\approx$50and $\approx$34in the smoke-positive prediction for the CNN and transformer models, respectively. These results underscore the need for further model training using data augmentation techniques. While transformers performed better during this test in our study, [14, 18] suggest that transformers are not necessarily stronger at patch perturbations than CNNs. Because this seems to be a continuity even for Big Data models, more research is recommended for this adversarial attack.
  • An analysis of the localization deception rate revealed areas within the images that were particularly vulnerable to local noise injection. Detailed analysis suggests that human annotation bias may cause to over-focus on the middle-horizontal region, offering insights for future data augmentation strategies to enhance model robustness. This is consistent with DNN behavior in other fields [10], highlighting the need to consider not only the visual characteristics of the target objects but also its spatiotemporal context for true unbiased training.
Based on our analysis, we propose the following data augmentation strategies for improving robustness. These solutions should diversify future training data for YOLOv8- and RT-DETR-based wildfire detection models. Moreover, because this framework is model-agnostic, these solutions may still be applied even if new DNN architectures are introduced.
@@
    • Gaussian noise with $0.1\lesssim a\lesssim 0.4$ should be introduced into training data, which should reduce precision degradation when encountering global noise, especially among speed-optimized transformers. It will also improve data variety and quantity.
@@
    • Clouds and smoke overlap most in the middle-horizontal strip of the images. Given that most false positives occur in this zone, cloud PNG patches should be placed in this zone to help models distinguish the two.
    • Furthermore, cloud PNG patches should also be placed into the upper areas (areas depicting the sky) to add spatial variety to local noise patches.
@@
    • An effective solution to combat false positives extensively used by [39] is to use collages. This is because collages allow models to easily compare between smoke-positive and smoke-negative images. Since CNNs suffered from local perturbation classification, collages of images between classes should be implemented into training data.
    • Furthermore, certain collage techniques such as the YOLOv4-style mosaic [4] has the added benefit of introducing object size variety in the data. This is particularly useful for small object detection.
    • However, collages inevitably make the already-small smoke object even smaller. Especially for speed-optimized transformers, which have a known weakness to small objects, crucial smoke features must be extracted first. The below augmentation strategy seeks to offer a potential solution.
@@
    • Data collected by [17] from [21] exclusively depict smoke that appears at or near the horizon (i.e., the middle-horizontal of the image). Cropping these images into equal quadrants creates four standalone images, which shift the position of objects. This adds spatial diversity to smoke annotations.
    • Furthermore, since the crops will result in a smaller-sized image, when resized to the target image size ($640\times 640$ pixels), the image, and by extension, smoke, will appear larger. This may help models better extract the subtle features of smoke, offering a solution to the problem discussed in the previous augmentation strategy.